OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -1/4x-5 y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (anonymous):

ok sinse it a'perp' you have to use the reciprocal for m. so your new slope is 4. then you use the formula \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

where \[y _{1}= -1 and x _{1}=7\]

OpenStudy (anonymous):

lol i have no clue what a perp is. and when you have y=4x how do you get the y?

OpenStudy (anonymous):

perp meand 90 degrees

OpenStudy (anonymous):

means*

OpenStudy (anonymous):

ohh okay then. do you have any clue how im post to find y?

OpenStudy (anonymous):

perpendicular and just plug the numbers using the formula and yu get your line

OpenStudy (anonymous):

oh thnks.

OpenStudy (anonymous):

well first use this equation y=mx+b where m is your slope and b is y intercept To find a slope perpendiculare you must must invert m and multiply by -1 your m is -1/4 so it is 4 if we invert and multiply by -1

OpenStudy (anonymous):

@hereweg0 The y value of the given point is 7, not -1.

OpenStudy (anonymous):

oh okai :/ i undeerstand everything to find the slope, but i still dnt see how to find the y intercept

OpenStudy (anonymous):

@rlongo yeah thanks. but just simplify and you get it too

OpenStudy (anonymous):

if you are looking for a line that passes through the point (-1,7) and you know the slope is 4, plug in the given numbers into the y intercept formula for a line. y=mx+b 7=4(-1)+b Solve for b

OpenStudy (anonymous):

well you are given Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -1/4x-5 and i showed you the formula y=mx+b and i said 4 is your slope so plug in (-1,7) 7=4*-1+b sfb b=11 so now y=4*x+11 :)

OpenStudy (anonymous):

oh i now understand it xD @timo86m this may sound mean but i didnt understand everyone elses explanation o.O lol

OpenStudy (anonymous):

Just curious, but what level of maths are you studying?

OpenStudy (anonymous):

im doing algebra 1 im doing florida virtual school, and they dont explain anything very well. But in August im starting geomotry in 10th grade

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Well because I am a great teacher :P

OpenStudy (anonymous):

haha okay

OpenStudy (anonymous):

I rushed btw lol

OpenStudy (anonymous):

and the mods warn me I suck at OS :P

OpenStudy (anonymous):

ah okay.

OpenStudy (anonymous):

they want me suspended :P

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

ur school???

OpenStudy (anonymous):

no this site the mods warn me

OpenStudy (anonymous):

lol whys that?

OpenStudy (anonymous):

lol cuzz i dont play by the rules. Or so they say.

OpenStudy (anonymous):

o.O theres rules?? haha

OpenStudy (anonymous):

there are a ton

OpenStudy (anonymous):

oh ops i hope i hvnt broken any hahaha xP

OpenStudy (anonymous):

i dont think you have lol. Basically so long as you not just looking for answers. They want you looking for explanations.

OpenStudy (anonymous):

oh really? thats stupid

OpenStudy (anonymous):

i think so too but i follow the rules now lol. SORTA

OpenStudy (anonymous):

lol 'Sorta' haha i just came on here to get awnsers for the questions i dont understand o.O

OpenStudy (anonymous):

they want you to learn not just look for answers.

OpenStudy (anonymous):

haha that's lame xP

OpenStudy (anonymous):

i agree. but what choice do we have lol?